### v1

@@ 3 Main Results
theoremsemiupper There exists an algorithm such that if there exists a graph $g^{\prime}\in\mathcal{H}$ with $\ell^{\textrm{precision}}(g^{\prime})=0$ and $\ell^{\textrm{recall}}(g^{\prime})=r$, then given an IID training set of size $O(\frac{\log(|\mathcal{H}|/\delta)}{\Delta_{\mathcal{D}}^{2}})$, with probability $1-\delta$, it outputs a graph with $\ell^{\textrm{precision}}(g^{\text{output}})=0$ and $\ell^{\textrm{recall}}(g^{\text{output}})=r$. When the target graph’s degree $n_{g^{\text{target}}}(x)$ is bounded by $C$ everywhere, we have
$\mathbb{E}\left[\frac{|N_{g}(x)\cap N_{g^{\text{target}}}(x)|}{n_{g}(x)\cdot n_{g^{\text{target}}}(x)}\right]=\mathbb{E}\left[\frac{1-\ell^{\textrm{precision}}(g,x)}{n_{g^{\text{target}}}(x)}\right]\leq\mathbb{E}\left[\frac{1}{n_{g^{\text{target}}}(x)}\right]-\mathbb{E}\left[\frac{\ell^{\textrm{precision}}(g,x)}{C}\right]\,.$
Therefore, we have $\Delta_{g,\mathcal{D}}\geq\frac{\ell^{\textrm{precision}}(g)}{C}$. This implies that when the target graph has bounded degree, we are able to find the graph with zero precision loss. However, when the target graph’s degree becomes too large, we show that it is impossible to achieve precision-recall of $(0,r)$.
@@
theoremsemilower There exists a class =$\mathcal{H}=\{g_{1},g_{2}\}$ of two graphs, for any $m>0$ and any (possibly randomized improper) algorithm $\mathcal{A}$, there exists a target graph $g^{\text{target}}$ and a data distribution $\mathcal{D}$ for which there exists a graph $g^{\dagger}\in\mathcal{H}$ with $\ell^{\textrm{precision}}(g^{\dagger})=0$ s.t. with probability $1-\delta$ over the training set, the expected (over the randomness of the algorithm) precision and recall losses of the output $g^{\text{output}}$ satisfy either $\mathbb{E}\left[\ell^{\textrm{recall}}(g^{\text{output}})\right]\geq\min_{g\in\mathcal{H}}\ell^{\textrm{recall}}(g)+\Omega(1)$ or $\mathbb{E}\left[\ell^{\textrm{precision}}(g^{\text{output}})\right]=\Omega(1)$. In the proof of the theorem, we construct a target graph with a degree much larger than the sample size $m$, as well as graphs $g_{1}$ and $g_{2}$ with degree $1$. In this setup, regardless of whether $g_{1}$ has perfect precision (i.e., the neighborhood in $g_{1}$ is a subset of the true neighborhood) or very poor precision, we cannot distinguish between these two cases because we never observe a neighbor in $g_{1}$ being sampled. Therefore, the only way to achieve zero precision loss is to output the empty graph, which, however, results in a high recall loss.
@@ 4 Proof Overview
Given a sequence of IID users $x_{1},\ldots,x_{m}$, let ^$\widehat{\ell}^{\textrm{precision}}(g)=\frac{1}{m}\sum_{i=1}^{m}\frac{|N_{g}(x_{i})\setminus N_{g^{\text{target}}}(x_{i})|}{n_{g}(x_{i})}$ and ^$\widehat{\ell}^{\textrm{recall}}(g)=\frac{1}{m}\sum_{i=1}^{m}\frac{|N_{g^{\text{target}}}(x_{i})\setminus N_{g}(x_{i})|}{n_{g^{\text{target}}}(x_{i})}$ denote the empirical precision and recall losses. It suffices to focus on empirical precision and recall losses minimization since by standard concentration bounds, minimizing these empirical losses leads to the minimization of the expected recall and precision losses.
@@
The maximum likelihood method returns $g^{\text{output}}=\operatorname*{arg\,max}_{g\in\mathcal{H}}\prod_{i=1}^{m}\frac{\mathds{1}(v_{i}\in N_{g}(x_{i}))}{n_{g}(x_{i})}$, which is equivalent to returning the graph with the minimum sum of log degrees among all consistent graphs, i.e.,
@@
Any consistent graph will have low empirical recall loss by applying standard concentration inequality and thus, ^$\widehat{\ell}^{\textrm{recall}}(g^{\text{output}})$ is small and we only to show that the empirical precision loss is small. The main technical challenge in the analysis is how to connect precision loss with likelihood.
@@
We first prove that for any graph $g$, its empirical precision loss can be bounded by a term of log degree and the empirical recall loss as follows:
^$\widehat{\ell}^{\textrm{precision}}(g)\leq\frac{2}{m}\sum_{i\in[m]:\frac{n_{g}(x_{i})}{n_{g^{\text{target}}}(x_{i})}\geq 1}\log\frac{n_{g}(x_{i})}{n_{g^{\text{target}}}(x_{i})}+2\widehat{\ell}^{\textrm{recall}}(g)\,.$ (2)
@@
^$\widehat{\ell}^{\textrm{precision}}(g^{\text{output}})\leq-\frac{2}{m}\sum_{i:\frac{n_{g^{\text{output}}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}<1}\log\frac{n_{g^{\text{output}}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}+2\widehat{\ell}^{\textrm{recall}}(g^{\text{output}})\,.$
However, with high probability, the first term $-\frac{2}{m}\sum_{i:\frac{n_{g^{\text{output}}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}<1}\log\frac{n_{g^{\text{output}}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}$ is small. This is because, for any graph $g$, the probability of outputting $g$ is
@@
For any graph $g$ with large $-\frac{2}{m}\sum_{i:\frac{n_{g}(x_{i})}{n_{g^{\text{target}}}(x_{i})}<1}\log\frac{n_{g}(x_{i})}{n_{g^{\text{target}}}(x_{i})}$, the probability of outputting such a graph is low.
@@
In the agnostic setting, all graphs in the hypothesis class may have zero likelihood of being the true graph; thus, the standard maximum likelihood method doesn’t work. However, we can make slight modifications to the maximum likelihood method to make it work for the Pareto-loss objective.
As mentioned earlier, the maximum likelihood method is equivalent to returning the graph with the minimum sum of log degrees among all consistent graphs. Hence, it can be decomposed into two steps: minimizing the recall loss and then regulating by minimizing the sum of log degrees. In the agnostic setting, given any $r$, we first find the set $\widehat{\mathcal{H}}$ of all graphs in the hypothesis class with recall loss at most $r+2\varepsilon$, and then regulate by minimizing the sum of log truncated degrees by returning
@@
where $a\wedge b:=\min(a,b)$. The truncation plays an important role here. Intuitively, let $g^{\dagger}$ denote the graph with precision and recall losses $p$ and $r$, respectively. If there exists an $x_{i}$ such that $n_{g^{\dagger}}(x_{i})$ is very large, minimizing the sum of log-untruncated degrees will never return $g^{\dagger}$. By applying truncation, we limit the effect of a single user with a very large degree.
@@
We formally capture this using a metric defined in the following. For any graph $g$, let $U_{i}^{g}$ denote the uniform distribution over the neighbors $N_{g}(x_{i})$ of $x_{i}$. Then, for any graph $g$, we define a vector $v_{g}:\mathcal{H}\times\mathcal{H}\to[0,1]$ by
$v_{g}(g^{\prime},g^{\prime\prime})=\frac{1}{m}\sum_{i=1}^{m}U_{i}^{g}(N_{g^{\prime}}(x_{i})\setminus N_{g^{\prime\prime}}(x_{i}))=\frac{1}{m}\sum_{i=1}^{m}\mathbb{P}_{v\sim U_{i}^{g}}\left(v\in N_{g^{\prime}}(x_{i})\setminus N_{g^{\prime\prime}}(x_{i})\right).$
Intuitively, $v_{g}(g^{\prime},g^{\prime\prime})$ represents the fraction of items users like that are recommended by $g^{\prime}$ but not by $g^{\prime\prime}$ in the counterfactual scenario where $g$ is the target graph. If $g$ is indeed the target graph, then this quantity should be consistent with our training data, i.e.,
$v_{g}(g^{\prime},g^{\prime\prime})\approx v_{\widehat{g}}(g^{\prime},g^{\prime\prime})\,,$
where $\widehat{g}$ is the observed (empirical) graph; i.e., the graph in which every $x_{i}$ is connected to the random number $v_{i}$ which is observed in the training set. In the above example, $v_{g^{\prime}}(g^{\prime},g^{\prime\prime})$ is 20% while $v_{\widehat{g}}(g^{\prime},g^{\prime\prime})$ is 5% and thus $g^{\prime}$ is unlikely to be the target graph.
We define a metric $d_{\mathcal{H}}$ between two graphs $g_{1}$ and $g_{2}$ by
@@
$\frac{1}{2}d_{\mathcal{H}}(g^{\text{target}},g)\leq\ell^{\textrm{scalar}}(g)\leq d_{\mathcal{H}}(g^{\text{target}},g)\,.$
@@
Then, we return a graph $g^{\text{output}}\in\mathcal{H}$ such that
@@
The Hardness of No Knowledge of the Target Graph’s Degree
For any graph $g$, it’s precision loss at any user $x$ is $\frac{|N_{g}(x)\setminus N_{g^{\text{target}}}(x)|}{n_{g}(x)}$ and we only get a random neighbor $v\sim N_{g^{\text{target}}}(x)$. If we are given the knowledge of the degree $n_{g^{\text{target}}}(x)$ of the target graph, then we can obtain an unbiased estimate of the precision loss, i.e., $1-\frac{n_{g^{\text{target}}}(x)}{n_{g}(x)}\cdot\mathds{1}(v\in N_{g}(x))$. But the difficulty lies in that we don’t know $n_{g^{\text{target}}}(x)$.
Consider the following example illustrated in Fig 3. For a given user $x$, there is a set of $n$ nodes equally divided into two sets $N_{1}(x)$ and $N_{2}(x)$. Consider two graphs—$g_{1}$ with neighborhood $N_{g_{1}}(x)=N_{1}(x)$ and $g_{2}$ with neighborhood $N_{g_{2}}(x)=N_{1}(x)\cup N_{2}(x)$ being all $n$ nodes.
In a world characterized by $\beta\in[\frac{1}{8},\frac{2}{3}]$, the target graph is generated in the following random way: Randomly select $\frac{3}{4}\cdot\beta n$ nodes from $N_{1}(x)$ and $\frac{1}{4}\cdot\beta n$ nodes from $N_{2}(x)$. No matter what $\beta$ is, w.p. $\frac{3}{4}$, $v$ is sampled uniformly at random from $N_{1}(x)$ and w.p. $\frac{1}{4}$, $v$ is sampled uniformly at random from $N_{2}(x)$. That is, every node in $N_{1}(x)$ has probability $\frac{3}{2n}$ of being sampled and every node in $N_{2}(x)$ has probability $\frac{1}{2n}$ of being sampled. Hence, if we have never seen the same user twice, we cannot distinguish between different $\beta$’s.
@@
$\ell^{\textrm{scalar}}(g_{1},x)=$ $1-\left(\frac{|N_{g^{\text{target}}}(x)\cap N_{g_{1}}(x)|}{2|N_{g_{1}}(x)|}+\frac{|N_{g^{\text{target}}}(x)\cap N_{g_{1}}(x)|}{2|N_{g^{\text{target}}}(x)|}\right)=1-(\frac{3/4\cdot\beta n}{n}+\frac{3/4\cdot\beta n}{2\beta n})$
@@
$\ell^{\textrm{scalar}}(g_{2},x)=1-\left(\frac{|N_{g^{\text{target}}}(x)\cap N_{g_{2}}(x)|}{2|N_{g_{2}}(x)|}+\frac{|N_{g^{\text{target}}}(x)\cap N_{g_{2}}(x)|}{2|N_{g^{\text{target}}}(x)|}\right)=1-\left(\frac{\beta n}{2n}+\frac{\beta n}{2\beta n}\right)=\frac{1}{2}-\frac{1}{2}\beta\,.$
When $\beta$ is large, $g_{1}$ has a smaller loss; when $\beta$ is small, $g_{2}$ has a smaller loss. With a huge number of users, we might never observe the same user twice and, therefore, cannot distinguish between different $\beta$ values. Consequently, it’s impossible to determine which of the two graphs has a smaller loss. We show that, in this example, even if algorithms are allowed to be randomized and improper, it still impossible to compete with the best graph in the hypothesis class.
$x$$N_{1}(x)$$N_{2}(x)$
(a) $g_{1}$
$x$$N_{1}(x)$$N_{2}(x)$
(b) $g_{2}$
$x$$N_{1}(x)$$N_{2}(x)$
(c) $g^{\text{target}}$ for small $\beta$
$x$$N_{1}(x)$$N_{2}(x)$
(d) $g^{\text{target}}$ for large $\beta$

### v2

@@ 3 Main Results
theoremsemiupper There exists an algorithm such that if there exists a hypothesis $g^{\prime}\in\mathcal{H}$ with $\ell^{\textrm{precision}}(g^{\prime})=0$ and $\ell^{\textrm{recall}}(g^{\prime})=r$, then given an IID training set of size $O(\frac{\log(|\mathcal{H}|/\delta)}{\Delta_{\mathcal{D}}^{2}})$, with probability $1-\delta$, it outputs a hypothesis with $\ell^{\textrm{precision}}(g^{\text{output}})=0$ and $\ell^{\textrm{recall}}(g^{\text{output}})=r$. When the target hypothesis’s output size $n_{g^{\text{target}}}(x)$ is bounded by $C$ almost everywhere, we have
$\mathbb{E}\left[\frac{|g(x)\cap g^{\text{target}}(x)|}{n_{g}(x)\cdot n_{g^{\text{target}}}(x)}\right]=\mathbb{E}\left[\frac{1-\ell^{\textrm{precision}}(g,x)}{n_{g^{\text{target}}}(x)}\right]\leq\mathbb{E}\left[\frac{1}{n_{g^{\text{target}}}(x)}\right]-\mathbb{E}\left[\frac{\ell^{\textrm{precision}}(g,x)}{C}\right]\,.$
Therefore, we have $\Delta_{g,\mathcal{D}}\geq\frac{\ell^{\textrm{precision}}(g)}{C}$. This implies that when the target hypothesis has bounded output size, we are able to find the hypothesis with zero precision loss. However, when the target hypothesis’s output size becomes too large, we show that it is impossible to achieve precision-recall of $(0,r)$.
@@
theoremsemilower There exists a class $\mathcal{H}=\{g_{1},g_{2}\}$ of two hypotheses, for any $m>0$ and any (possibly randomized improper) algorithm $\mathcal{A}$, there exists a target hypothesis $g^{\text{target}}$ and a data distribution $\mathcal{D}$ for which there exists a hypothesis $g^{\dagger}\in\mathcal{H}$ with $\ell^{\textrm{precision}}(g^{\dagger})=0$ s.t. with probability $1-\delta$ over the training set, the expected (over the randomness of the algorithm) precision and recall losses of the output $g^{\text{output}}$ satisfy either $\mathbb{E}\left[\ell^{\textrm{recall}}(g^{\text{output}})\right]\geq\min_{g\in\mathcal{H}}\ell^{\textrm{recall}}(g)+\Omega(1)$ or $\mathbb{E}\left[\ell^{\textrm{precision}}(g^{\text{output}})\right]=\Omega(1)$. In the proof of the theorem, we construct a target hypothesis with output size much larger than the sample size $m$, as well as hypotheses $g_{1}$ and $g_{2}$ with output size $1$. In this setup, regardless of whether $g_{1}$ has perfect precision (i.e., the output of $g_{1}$ is a subset of the ground-truth set) or very poor precision, we cannot distinguish between these two cases because we never observe a label in the output of $g_{1}$ being sampled. Therefore, the only way to achieve zero precision loss is to output the empty set function, which, however, results in a high recall loss.
@@ 4 Proof Overview
Given a sequence of IID inputs $x_{1},\ldots,x_{m}$, let
$\widehat{\ell}^{\textrm{precision}}(g)=\frac{1}{m}\sum_{i=1}^{m}\frac{|g(x_{i})\setminus g^{\text{target}}(x_{i})|}{n_{g}(x_{i})}$
and
$\widehat{\ell}^{\textrm{recall}}(g)=\frac{1}{m}\sum_{i=1}^{m}\frac{|g^{\text{target}}(x_{i})\setminus g(x_{i})|}{n_{g^{\text{target}}}(x_{i})}$
denote the empirical precision and recall losses. It suffices to focus on empirical precision and recall losses minimization since by standard concentration bounds, minimizing these empirical losses leads to the minimization of the expected recall and precision losses.
@@
The maximum likelihood method returns $g^{\text{output}}=\operatorname*{arg\,max}_{g\in\mathcal{H}}\prod_{i=1}^{m}\frac{\mathds{1}(v_{i}\in g(x_{i}))}{n_{g}(x_{i})}$, which is equivalent to returning the hypothesis with the minimum sum of log output sizes among all consistent hypotheses, i.e.,
@@
Any consistent hypothesis will have low empirical recall loss by applying standard concentration inequality and thus, $\widehat{\ell}^{\textrm{recall}}(g^{\text{output}})$ is small and we only need to show that the empirical precision loss is small. The main technical challenge in the analysis is how to connect precision loss with likelihood.
@@
We first prove that for any hypothesis $g$, its empirical precision loss can be bounded by a term of log output size and the empirical recall loss as follows:
$\widehat{\ell}^{\textrm{precision}}(g)\leq\frac{2}{m}\sum_{i\in[m]:\frac{n_{g}(x_{i})}{n_{g^{\text{target}}}(x_{i})}\geq 1}\log\frac{n_{g}(x_{i})}{n_{g^{\text{target}}}(x_{i})}+2\widehat{\ell}^{\textrm{recall}}(g)\,.$ (2)
@@
$\widehat{\ell}^{\textrm{precision}}(g^{\text{output}})\leq-\frac{2}{m}\sum_{i:\frac{n_{g^{\text{output}}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}<1}\log\frac{n_{g^{\text{output}}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}+2\widehat{\ell}^{\textrm{recall}}(g^{\text{output}})\,.$
However, with high probability, the first term $\frac{2}{m}\sum_{i:\frac{n_{g^{\text{output}}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}<1}\log\frac{n_{g^{\text{output}}}(x_{i})}{n_{g^{\text{target}}}(x_{i})}$ is small. This is because, for any hypothesis $g$, the probability of outputting $g$ is
@@
For any hypothesis $g$ with large $\frac{2}{m}\sum_{i:\frac{n_{g}(x_{i})}{n_{g^{\text{target}}}(x_{i})}<1}\log\frac{n_{g}(x_{i})}{n_{g^{\text{target}}}(x_{i})}$, the probability of outputting such a hypothesis is low.
@@
In the agnostic setting, all hypotheses in the hypothesis class may have zero likelihood of being the true hypothesis; thus, the standard maximum likelihood method doesn’t work. However, we can make slight modifications to the maximum likelihood method to make it work for the Pareto-loss objective.
As mentioned earlier, the maximum likelihood method is equivalent to returning the hypothesis with the minimum sum of log output sizes among all consistent hypotheses. Hence, it can be decomposed into two steps: minimizing the recall loss and then regulating by minimizing the sum of log output sizes. In the agnostic setting, given any $r$, we first find the set $\widehat{\mathcal{H}}$ of all hypotheses in the hypothesis class with recall loss at most $r+2\varepsilon$, and then regulate by minimizing the sum of log truncated output sizes by returning
@@
where $a\wedge b:=\min(a,b)$. The truncation plays an important role here. Intuitively, let $g^{\dagger}$ denote the hypothesis with precision and recall losses $p$ and $r$, respectively. If there exists an $x_{i}$ such that $n_{g^{\dagger}}(x_{i})$ is very large, minimizing the sum of log-untruncated output sizes will never return $g^{\dagger}$. By applying truncation, we limit the effect of a single input with a very large output size.
@@
We formally capture this using a metric defined in the following. For any hypothesis $g$, let $U_{i}^{g}$ denote the uniform distribution over the output set $g(x_{i})$ at $x_{i}$. Then, for any hypothesis $g$, we define a vector $v_{g}:\mathcal{H}\times\mathcal{H}\to[0,1]$ by
$v_{g}(g^{\prime},g^{\prime\prime})=\frac{1}{m}\sum_{i=1}^{m}U_{i}^{g}(g^{\prime}(x_{i})\setminus g^{\prime\prime}(x_{i})):=\frac{1}{m}\sum_{i=1}^{m}\mathbb{P}_{v\sim U_{i}^{g}}\left(v\in g^{\prime}(x_{i})\setminus g^{\prime\prime}(x_{i})\right).$
Intuitively, $v_{g}(g^{\prime},g^{\prime\prime})$ represents the fraction of items inputs like that are recommended by $g^{\prime}$ but not by $g^{\prime\prime}$ in the counterfactual scenario where $g$ is the target hypothesis. If $g$ is indeed the target hypothesis, then this quantity should be consistent with our training data, $v_{g}(g^{\prime},g^{\prime\prime})\approx v_{\widehat{g}}(g^{\prime},g^{\prime\prime})$, where $\widehat{g}$ is the observed (empirical) hypothesis; i.e., the hypothesis in which every $x_{i}$ is mapped to the random number $v_{i}$ which is observed in the training set. In the above example, $v_{g^{\prime}}(g^{\prime},g^{\prime\prime})$ is 20% while $v_{\widehat{g}}(g^{\prime},g^{\prime\prime})$ is 5% and thus $g^{\prime}$ is unlikely to be the target hypothesis.
We define a metric $d_{\mathcal{H}}$ between two hypotheses $g_{1}$ and $g_{2}$ by
@@
$d_{\mathcal{H}}(g^{\text{target}},g)\leq 2\ell^{\textrm{scalar}}(g)\leq 4d_{\mathcal{H}}(g^{\text{target}},g)+2\min_{g^{\prime}\in\mathcal{H}}\ell^{\textrm{scalar}}(g^{\prime})\,.$
@@
Then, we return a hypothesis $g^{\text{output}}\in\mathcal{H}$ such that
@@
The Hardness of No Knowledge of the Target Hypothesis’s Output Size
For any hypothesis $g$, it’s precision loss at any input $x$ is $\frac{|g(x)\setminus g^{\text{target}}(x)|}{n_{g}(x)}$ and we only get a random label $v\sim\mathrm{Unif}(g^{\text{target}}(x))$. If we are given the knowledge of the output size $n_{g^{\text{target}}}(x)$ of the target hypothesis, then we can obtain an unbiased estimate of the precision loss, i.e., $1-\frac{n_{g^{\text{target}}}(x)}{n_{g}(x)}\cdot\mathds{1}(v\in g(x))$. But the difficulty lies in that we don’t know $n_{g^{\text{target}}}(x)$.
[…]
In a world characterized by $\beta\in[\frac{1}{8},\frac{2}{3}]$, the target hypothesis is generated in the following random way: Randomly select $\frac{3}{4}\cdot\beta n$ items from $N_{1}(x)$ and $\frac{1}{4}\cdot\beta n$ items from $N_{2}(x)$. No matter what $\beta$ is, w.p. $\frac{3}{4}$, $v$ is sampled uniformly at random from $N_{1}(x)$ and w.p. $\frac{1}{4}$, $v$ is sampled uniformly at random from $N_{2}(x)$. That is, every item in $N_{1}(x)$ has probability $\frac{3}{2n}$ of being sampled and every item in $N_{2}(x)$ has probability $\frac{1}{2n}$ of being sampled. Hence, if we have never seen the same input twice, we cannot distinguish between different $\beta$’s.
@@
$\ell^{\textrm{scalar}}(g_{1},x)=$ $1-\left(\frac{|g^{\text{target}}(x)\cap g_{1}(x)|}{2|g_{1}(x)|}+\frac{|g^{\text{target}}(x)\cap g_{1}(x)|}{2|g^{\text{target}}(x)|}\right)=1-(\frac{3/4\cdot\beta n}{n}+\frac{3/4\cdot\beta n}{2\beta n})$
@@
$\ell^{\textrm{scalar}}(g_{2},x)=1-\left(\frac{|g^{\text{target}}(x)\cap g_{2}(x)|}{2|g_{2}(x)|}+\frac{|g^{\text{target}}(x)\cap g_{2}(x)|}{2|g^{\text{target}}(x)|}\right)=1-\left(\frac{\beta n}{2n}+\frac{\beta n}{2\beta n}\right)=\frac{1}{2}-\frac{1}{2}\beta\,.$
When $\beta$ is large, $g_{1}$ has a smaller loss; when $\beta$ is small, $g_{2}$ has a smaller loss. With a huge input space $\mathcal{X}$ and the distribution $\mathcal{D}$ being uniform over such a huge space $\mathcal{X}$, it is almost sure that we will never observe the same input twice and, therefore, cannot distinguish between different $\beta$ values. Consequently, it’s impossible to determine which of the two hypotheses has a smaller loss. We show that, in this example, even if algorithms are allowed to be randomized and improper, it still impossible to compete with the best hypothesis in the hypothesis class.
[…]
Figure 3(a): $g_{1}$
[…]
Figure 3(b): $g_{2}$
[…]
Figure 3(c): $g^{\text{target}}$ for small $\beta$
[…]
Figure 3(d): $g^{\text{target}}$ for large $\beta$